Circular No. 6813

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

VARIABLE OBJECT IN DORADO D. L. Welch, McMaster University, reports that the MACHO collabo-

ration (cf. IAUC 6312) has discovered what appears to be a variable star in images obtained during monitoring of the Large Magellanic Cloud. The object peaked in brightness around 1997 Dec. 22 UT at V = 19.0 and is located at α = 4^{h}57^{m}46^{s}.2, δ = −67^{o}41′08′′ (equinox 2000.0). The last image before outburst was taken on Dec. 3.511, and the next observation, on Dec. 14.494, indicates that the object had brightened by at least 1.3 mag to V ∼ 19.1. Eight pairs of red and blue images taken after outburst are consistent with typical supernova fading timescales, and there is a possible galaxy located 3′′ to the west of the variable (finder charts can be obtained

at http://wwwmacho.mcmaster.ca/SN/index.html). It is therefore pos-

sible that this variable is a supernova in a galaxy beyond the LMC. No activity is present in 160 prior observations stretching back two years.

MXB 1730–335 R. Rutledge, University of California at Berkeley; C. Moore, Kapteyn

Institute; D. Fox and W. Lewin, Massachusetts Institute of Technology; and J. van Paradijs, University of Amsterdam and University of Alabama at Huntsville, report the discovery of a radio transient with flux density correlated with the RXTE/ASM x-ray flux of the rapid burster MXB 1730–335 (cf. IAUC 6409, 6506), in five observations at the Very Large Array at 8.0 GHz during the Dec. 1996 and June/July 1997 outbursts. The position of the radio source is α = 17^{h}33^{m}24^{s}.61, δ = −33^{o}23′19′′.8 (equinox 2000.0; uncertainty 0′′.1). Details for a proposed radio campaign during the next outburst (expected in the next few months) will be given

at http://astron.berkeley.edu/~rutledge/rb.html.

COMET C/1997 S2 (SOHO) Further to IAUC 6811, another comet, tail-less and with peak mag

∼ 7.5, has been found by D. A. Biesecker in the SOHO-LASCO archive, bringing the total number of SOHO comets to 40. Details are on MPEC 1998-B22.

1997 UT  α_{2000}  δ_{2000}
Sept.25.442  11^{h}44.^{m}9  − 2^{o}15′
1998 January 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green